OpenStudy (anonymous):

(1,-1), (-4, 5) I'm trying to find the equation of the line that passes through the points

OpenStudy (anonymous):

find the slope first which is y2 - y1divided by x2 - x1, and than put it in slope form y=mx+b.

OpenStudy (anonymous):

slope would be -.5

OpenStudy (anonymous):

5-(-1) = 5+ 1 = 6 and then over -4 - 1 = -5 so the slope would be -6/5

OpenStudy (anonymous):

m = (5-(-1))/(-4-1) so m= .5

OpenStudy (anonymous):

then take any point and plug it into y = mx + b and solve for b. so take (1,-1) and plug it in -1 = (-6/5)(1) + b -1 = -6/5 + b b = -1 + 6/5 = 1/5

OpenStudy (anonymous):

what's 5 - (-1)?

OpenStudy (anonymous):

it's 6..., and -4 -1 is -5

OpenStudy (anonymous):

that negative makes the negative in the bracket a positive.

OpenStudy (anonymous):

so your finished equation will be y = -6/5x + 1/5

OpenStudy (anonymous):

Oh I wrote it down without the changed sign.